the larger of two consecutive even integers is 2 more than twice the smaller. find the numbers.
Answer by checkley77(12569)

Let x & (x+2) be the 2 even integers.
x+2=2x+2
x-2x=2-2
-x=0
x=0 ans. for the smaller integer.
x+2=2 ans. for the larger integer.
Proof:
2=2*0+2
2=2